OpenStudy (anonymous):

What is the value for x in the proportion x over x minus six equals four over seven?

OpenStudy (anonymous):

you cross multiply: x(7)= x-6(4) and solve for x

OpenStudy (anonymous):

x/x-6 = 4/7 Cross multiply 7x = 4(x-6) 7x = 4x - 24 3x = -24 x =-8

OpenStudy (anonymous):

\[\frac{x}{x-6}=\frac{4}{7}\] 7x=4x-24 3x=-24 x=-8